# Dynamical phases of short-term memory mechanisms in RNNs

### v1

@@ Abstract
Short-term memory is essential for cognitive processing, yet our understanding of its neural mechanisms remains unclear. A key focus in neuroscience has been the study of sequential activity patterns, where neurons fire one after another within large networks, to explain how information is maintained. While recurrent connections were shown to drive sequential dynamics, a mechanistic understanding of this process still remains unknown. In this work, we first introduce two unique mechanisms that can subserve short-term memory: slow-point manifolds generating direct sequences or limit cycles providing temporally localized approximations. Then, through analytical models, we identify fundamental properties that govern the selection of these mechanisms, i.e., we derive theoretical scaling laws for critical learning rates as a function of the delay period length, beyond which no learning is possible. We empirically verify these observations by training and evaluating more than 35,000 recurrent neural networks (RNNs) that we will publicly release upon publication. Overall, our work provides new insights into short-term memory mechanisms and proposes experimentally testable predictions for systems neuroscience.
@@ 1 Introduction
Short-term memory is the process through which both biological and artificial systems are able to store information for a short period of time. This process is essential for adaptive behavior, which is disrupted in several neurological and psychiatric conditions such as Alzheimer’s disease [1], schizophrenia [2] and post-traumatic stress disorder [3] among others. Yet, while short-term memory is a central function in mammalian cognition, the mechanisms that subserve it remain an open question in systems neuroscience [4].
In both humans and animals, short-term memory has been shown to be supported through a distributed dynamic network across cortical regions such as prefrontal, parietal and sensory regions [5, 6, 7, 8, 9, 10, 11]. At the neuronal level, persistent neuronal activity has been linked to the memory maintenance period [8, 12, 6, 13] (Fig. 1). Some other evidence showed that short-term memory can also be supported in the form of sequential neuronal activations [9, 14, 15, 16] (Fig. 1). However, growing evidence suggests distributed and dynamic firing patterns across the neuronal population [10, 17, 11], where memory stability and robustness is achieved through attractor dynamics [18, 19]. It is unclear what are the different factors influencing how memory is maintained in the neuronal activities, what is the underlying mechanism, and how it changes over extended delay periods.
Here we set out to answer these questions through a computational theory perspective, which in turn provide experimental testable predictions. To achieve that, we trained over 35,000 RNNs to perform delayed cue discrimination tasks, while studied simple toy models and low-rank RNN models as they performed a simplified delayed activation task. In our tasks, neural activation sequences emerged as the precursor of the solutions. These sequences were generated through two distinct mechanisms: slow-point manifolds and limit cycles. Surprisingly, limit cycles emerged naturally to approximate neuronal sequences, even though the task had no periodicity component, offering a novel mechanism on how these sequences may be generated in the brain. An analytical study revealed that the selection of the mechanism (i.e. slow-point manifold or limit cycles) responsible for the sequences depended on both the task structure and the learning rate of the RNNs. Overall, our results provide new insights and testable experimental predictions for systems neuroscience. We summarize our unique contributions in this paper below:
  • Using interpretable dynamical system models stripped down to their most essential components for solving a delayed activation task, we illustrate how RNNs can develop slow-point manifold or limit cycle strategies during learning, depending on the task and optimization parameters (Figs. 2 and 3).
  • We show how changing one of the simplest tasks in a trivial manner, e.g., by adding a post-reaction period, can qualitatively change the learned short-term memory mechanisms. This finding is particularly relevant as a cautionary tale for the systems neuroscience community modeling animal behaviors by drawing comparisons to RNNs [20] (Fig. 4).
  • Utilizing analytical models, we demonstrate theoretical scaling laws and synthesize a phase diagram of mechanisms for solving the task, which is a function of the task properties and optimization parameters (Fig. 5).
  • Through a large-scale study of over 35,000 RNNs (which will be made publicly available), we extract scaling laws in RNNs trained to perform delayed cue-discrimination tasks. We show that RNNs have little to no trouble learning a modified task with no delayed output, providing further evidence that training neural networks to delay their outputs is a fundamentally difficult problem (Figs. 6, 7, S1, and S2).
Figure 1: A visualization of existing prevalent theories of working memory maintenance in neural activations. Left. Earlier models predict that fixed and persistent activities of a subset of neurons during the delay period could encode the memory content [21, 22]. Middle. Memory content can also be stored within sequentially firing neuron groups [15]. Right. Limit cycles can also play a significant role in working memory maintenance. However, if left undisturbed, they would continue to provide a response in the post-reaction window (magenta).
@@ 2 Background
Systems neuroscience often utilizes animal models, e.g., mice, rats, birds, flies, to study cognitive functions including short-term memory [4]. Similarly, it is often possible to design artificial models such as RNNs, which can be trained on observed neural activities or tasks that are traditionally performed by animals in laboratories [20, 23, 24, 25, 26]. Then, once RNNs are trained, their computation can be reverse-engineered to draw potential inferences to the nature of the neural computation performed by biological networks that can later be tested with experimental procedures [27, 28]. Consequently, RNNs have recently become a key component of computational neuroscience studies [26, 29, 28, 23].
Due to their biological relevance and easy interpretation of their parameters as functional connections between neurons [30, 31], in this work, we study RNNs with the following set of equations:
@@
where $r(t)\in\mathbb{R}^{N}$ refers to the firing rates, $u(t)\in\mathbb{R}^{N_{\rm in}}$ pre-defined inputs, $W/W^{\rm in}$ weight parameters and $b$ corresponding bias terms, and $\epsilon$ some noise that is, in practice, sampled iid from a Gaussian distribution. The output is taken as a projection of the neural activities as $\hat{o}(t)=f\left(W^{\rm out}r(t)+b_{\rm out}\right)$, where $f(\cdot)$ is either identity or sigmoid function in this work (see below). For notational simplicity, we assume $\epsilon=0$ until our empirical results in Section 3.4.
Recent work has shown that when $W=\sum_{k=1}^{K}m^{(k)}n^{(k)T}$ are constrained to be low-rank, the computation performed by RNNs can be written in terms of a latent dynamical system [26, 29, 32]:
$\tau\dot{\kappa}_{k}(t)=-\kappa_{k}(t)+\sum_{i=1}^{N}n_{i}^{(k)}\tanh\left(\sum_{l=1}^{K}m^{(l)}\kappa_{l}(t)+W^{\rm in}u(t)+b\right),$ (2)
where the latent variables are defined as $\kappa_{k}(t)=n^{(k)T}r(t)$ [32]. Intuitively, by constraining the rank of the weight matrix, $W$, one can project the computation performed by $N$ neurons down to $K$-dimensional latent dynamical systems. When we study low-rank RNNs, we define the network output as a nonlinear projection from the latent variables with a sigmoid function $f(\cdot)$. In this work, we use this paradigm (i.e., training low-rank RNNs) to illustrate the distinct mechanisms that can emerge during training to support sequence generation. Then, our large-scale experiments generalize our findings to full-rank RNNs with no explicit restrictions.
@@ 3 Results
Using a slightly modified version111which is nonetheless computationally equivalent after a simple transformation of variables [33]. of the architecture in Eq. (1), recent work has identified generation of neural sequences as a potential key component of short-term memory [15]. This sequence then can be used to store the short-term memory content, while allowing to explicitly keep track of the time, as opposed to earlier theories of short-term memory maintenance that relied on persistent activities of some neural populations [21, 22] (Fig. 1). Previous work has provided insights into how these sequences are generated by examining neuronal-level mechanisms [15, 34, 35], often focusing on the learned connectivity structures [15, 35]. However, the causal mechanisms underlying neural computation, which could be investigated through latent dynamics, remain largely unexplored.
3.1 RNNs can generate neural sequences using distinct mechanisms
@@
As a first step, we set out to design a simple short-term memory task, where any extra component is stripped away, that can be solved using neural sequences but not persistent firings of the same groups of neurons (Fig. 1). Inspired by the delayed cue discrimination tasks regularly studied in systems neuroscience [36], we started by considering a delayed activation task [37]. In this task, the network is initialized to a particular state, which is randomly selected in the state space (Methods), and has to withhold its response ($\hat{o}(t)=0$) for a $T_{\rm delay}$ time window. Afterwards, the network is forced to output a response ($\hat{o}(t)=1$) for $T_{\rm resp}$ time window.
@@
Notably, such a task requires at least two distinct neural activity patterns in two periods and therefore cannot be solved by persistent activity of a subset of neurons. Instead, when we trained rank-2 RNNs to solve the delayed activation task, we observed the emergence of neural sequences (Fig. 2). Interestingly, when we studied the flow maps of the learned dynamical systems, we observed that both limit cycles and a mechanism utilizing slow-points (from now on referred to as “SP manifolds” for generality) could generate the desired sequence in the trial window. Despite having equivalent results within the trial window, the two mechanisms had qualitatively different behaviors after the trial concluded (Fig. 2). Limit cycles had led to recurrent generations of the same sequences, whereas the SP manifolds converged to a persistent activity style representation. Here, we observed that, for a given delay interval $T_{\rm delay}$, networks tended to form limit cycles when trained with large learning rates, but could form SP manifolds in lower learning rates (Fig. 2). We will make this observation rigorous in our large-scale studies below, but for now, we asked the question: Do RNNs commit to learning one mechanism from the start, or are they capable of changing their inner mechanisms during the training?
@@
To answer this question, we studied the latent dynamical systems of a rank-2 RNN as it was solving the delayed activation task (Fig. 3). Interestingly, the loss function showed characteristic periods of learning, corresponding to changes in mechanisms (Fig. 3A). Specifically, after the first jump in the loss function, a SP manifold was formed in the latent dynamical system (Fig. 3B; left). Here, a line of slow-points allowed the network output to stall until $T_{\rm delay}$. Then, the SP manifold had acquired a curved structure (Fig. 3B; middle), reminiscent of a transitory solution. Finally, after the oscillations in the loss function values ceased, a limit cycle had emerged in the latent dynamical system (Fig. 3B; right). Hence, we confirmed that RNNs could change their inner mechanisms during learning, i.e., what started as a SP manifold solution later evolved into a limit cycle.
@@
In reality, since functional connections in the brain can rapidly reorganize within the start and end of a trial [36], the two mechanisms can become practically inseparable in biological networks. In other words, there is an equivalence class of mechanisms that can generate neural sequences that have the same local behavior within the task window, but can show distinct properties outside; and the two mechanisms we identified belong to this class. For the rest of this work, we aim to understand whether these mechanisms emerge randomly, or whether there is some structure to which mechanism is learned by the network based on task and optimization parameters.
Figure 2: RNNs use diverse mechanisms to solve the delayed activation task. To illustrate that different mechanisms can be employed by dynamical systems to delay their responses, we trained a rank-2 RNN to perform the delayed activation task. In this task, networks initialized at a given state (Methods) should delay their output by $T_{\rm delay}$ for a $T_{\rm resp}$ time interval. The task does not constrain the network activity or the output outside of these time windows. A Left. The normalized gradient norm and the loss function values. Middle. We probe the learned latent mechanisms by plotting the flow maps of the latent dynamical systems. The shaded areas around the decision boundary correspond to network outputs within $0.75$ (outer green boundary) and $0.25$ (outer red boundary). Right. We plotted the (ordered) firing rates for the neurons belonging to the networks illustrated in the middle column, both during and after the trial. Parameters: $N=100$ neurons, $\tau=10$ms, $\Delta t=5$ms, $T_{\rm delay}=75$ms, $T_{\rm resp}=50$ms, $\alpha=1$ for stochastic gradient descent using otherwise the default parameters in Pytorch [38]. B-E Same as in panel (A), but with varying learning rates. For the largest learning rate (A), RNN was incapable of solving the task despite the substantial vanishing of the gradient. As the learning rates decreased, the RNN was capable of solving the task with limit cycles in moderate to high values of the learning rates (B-C), but utilized slow-points for the lowest learning rates (D-E). Both mechanisms were capable of producing neural activity sequences during the task window.
Figure 3: An overview of the emergent mechanisms in a rank-2 RNN solving the delayed activation task. A To illustrate that different mechanisms can be employed by dynamical systems to delay their responses, we trained a rank-2 RNN to perform the delayed activation task. In this task, networks initialized at a given state (Methods) should delay their output by $T_{\rm delay}$. Parameters: $N=100$ neurons, $\tau=10$ms, $\Delta t=5$ms, $T_{\rm delay}=150$ms, $T_{\rm resp}=50$ms, $\alpha=10^{-2}$ for stochastic gradient descent using otherwise the default parameters in Pytorch [38]. The dotted lines correspond to epochs, for which the latent dynamical systems are illustrated in Panel (B). B During the training, two distinct mechanisms emerge in the RNN, which we probe by plotting the two-dimensional latent dynamical systems. Left. Early in the training, the network lowers the loss values by designing a line of slow-points (referred to as SP manifold), towards which other stats are attracted. The latent states after the trial concludes end up in a final attractive fixed-point. Middle. As training progresses, the SP manifold slowly dissolves into a circular pattern, which we deem as a transient mechanism before the network settles into the final solution. Right. After further training, a limit cycle emerges in the latent dynamical system. In this structure, even after the trial concludes, latent states continue to evolve in a periodic manner. The shaded areas around the decision boundary correspond to network outputs within $0.75$ (outer green boundary) and $0.25$ (outer red boundary).
3.2 Influence of task design on the learned latent mechanism

### v2

@@ Abstract
Short-term memory is essential for cognitive processing, yet our understanding of its neural mechanisms remains unclear. Neuroscience has long focused on how sequential activity patterns, where neurons fire one after another within large networks, can explain how information is maintained. While recurrent connections were shown to drive sequential dynamics, a mechanistic understanding of this process still remains unknown. In this work, we introduce two unique mechanisms that can support this form of short-term memory: slow-point manifolds generating direct sequences or limit cycles providing temporally localized approximations. Using analytical models, we identify fundamental properties that govern the selection of each mechanism. Precisely, on short-term memory tasks (delayed cue-discrimination tasks), we derive theoretical scaling laws for critical learning rates as a function of the delay period length, beyond which no learning is possible. We empirically verify these results by training and evaluating approximately 80,000 recurrent neural networks (RNNs), which are publicly available for further analysis [^1] https://github.com/fatihdinc/dynamical-phases-stm. Overall, our work provides new insights into short-term memory mechanisms and proposes experimentally testable predictions for systems neuroscience.
Neuroscience, recurrent neural networks
@@ 1 Introduction
Learning to respond based on delayed information is a fundamental challenge in both neuroscience and machine learning (Constantinidis & Klingberg, 2016; Bengio et al., 1994). In machine learning, early studies in the 1990s identified this difficulty as the problem of learning long-term dependencies, particularly in recurrent neural networks (RNNs) trained on tasks requiring information retention across time (Bengio et al., 1994). Despite significant progress in neural network architectures, this issue remains central to designing systems capable of effective temporal reasoning.
Figure 1: Studying short-term memory mechanisms with biologically motivated neural network models. A-C A visualization of short-term memory maintenance in neural activations. A Earlier models predict that fixed and persistent activities of a subset of neurons during the delay period could encode the content corresponding to distinct memories (red vs blue populations) (Fuster & Alexander, 1971; Atkinson & Shiffrin, 1968). B Memory contents can also be stored within a sequentially firing neuron population (Rajan et al., 2016), where distinct readout neurons decode distinct memories. C In this work, we show that limit cycles, i.e., periodic orbits, can locally approximate the memory mechanism in panel (B). However, if left undisturbed, limit cycles may result in spurious periodic responses in the post-reaction window (magenta). D-E An illustration of the delayed cue-discrimination task and its simulation with RNNs. D The task includes two correct cue-reaction matches (blue-left, red-right) and consists of four periods: (i) Cue period, where one of the two cues is presented; (ii) Delay period, with no expected reaction; (iii) Reaction period, where the cue class needs to be outputted; and (iv) an optional post-reaction period at the end of the trial. E We designed an equivalent task for RNNs, in which two input channels are mapped to the output channels and for a given trial only a single input is provided in the form of a rectangular pulse of duration $T_{\rm in}$. RNNs can be trained to follow the stages (i-iv) similar to the animals in their target responses. When the cue period is omitted, we refer to this task as delayed activation task.
To address long-term dependencies, machine learning researchers developed specialized architectures such as Long Short-Term Memory (LSTM) networks, Gated Recurrent Units (GRUs), and, more recently, Transformers (Hochreiter, 1997; Chung et al., 2014; Vaswani, 2017). These models introduced mechanisms that enhance memory retention, e.g., by learning neuronal timescales. In contrast, biological neurons communicate through action potentials and are governed by biophysical parameters, such as membrane time constants, that determine the decay of activity over time (Jonas et al., 1993; McCormick et al., 1985; Miles, 1990; Isokawa, 1997; Geiger et al., 1997). This fundamental difference raises a key question: how do biologically inspired neural systems learn to associate events separated in time?
@@
This question has regularly been studied under a central topic in neuroscience: short-term memory, a fundamental cognitive function that enables organisms to temporarily store and manipulate information. This process is essential for adaptive behavior and is disrupted in several neurological and psychiatric conditions such as Alzheimer’s disease (Bondi et al., 2017), schizophrenia (Lee & Park, 2005) and post-traumatic stress disorder (Scott et al., 2015) among others. Yet, while short-term memory is a central cognitive function, the mechanisms that support it remain poorly understood in systems neuroscience (Serences, 2016).
In the mammalian brain, short-term memory is supported by a dynamic network spanning cortical and subcortical regions such as the prefrontal, parietal, and sensory cortices, and the hippocampus (Todd & Marois, 2004; Curtis & D’Esposito, 2003; Harrison & Tong, 2009; Gnadt & Andersen, 1988; Baeg et al., 2003; Cavanagh et al., 2018; Meyers et al., 2008). Theories explaining this core process can be broadly grouped into three levels. At the synaptic transmission level, memory may be stored in “activity-silent” states, maintained by transient synaptic changes induced by plasticity (Mongillo et al., 2008; Stokes, 2015; Kozachkov et al., 2022; Masse et al., 2019; Brennan & Proekt, 2023). At the neuronal activity level, persistent spiking has been linked to memory maintenance during delay periods (Gnadt & Andersen, 1988; Barash et al., 1991; Curtis & D’Esposito, 2003; Baddeley & Hitch, 1974) (Fig. 1A). Finally, at the population level, short-term memory has been associated with sequential patterns of neural activation (Baeg et al., 2003; Harvey et al., 2012; Rajan et al., 2016; Wang, 2021) (Fig. 1B).
In this work, we study mechanisms that can support memory maintenance at the final population level (Fig. 1B-C). Increasing evidence suggests that memory stability and robustness can emerge from a dynamic, distributed neural population activity and is governed by underlying attractor states (Cavanagh et al., 2018; Spaak et al., 2017; Meyers et al., 2008; Stroud et al., 2024; Brennan & Proekt, 2023). Yet several theoretical, and empirically relevant, questions remain unresolved: Q1: What mechanisms support memory maintenance through sequential neural activity? Q2: What determines which mechanism is favored under different memory tasks or training conditions? Q3: And how do these learned mechanisms vary with the delay duration, i.e., the time during which the information must be held in memory?
We set out to answer these questions through a computational theory perspective, generating predictions for in-vivo experiments that we then test in-silico. To achieve this, we trained approximately 80,000 recurrent neural networks (RNNs) to perform classical system neuroscience experiments designed to assess short-term memory capabilities, while also studying simple toy models and low-rank RNN models as they performed a simpler short-term memory task (Fig. 1D-E). Our contributions can be summarized as follows: (i) Using interpretable dynamical system models stripped down to their most essential components for solving a delayed activation task (Fig. 1D-E), we illustrate how RNNs can develop distinct strategies during learning, depending on the task and optimization parameters (Figs. 2, 3, and S1). (ii) We show how changing one of the simplest tasks in a trivial manner, e.g., by adding a post-reaction period, can qualitatively change the learned short-term memory mechanisms. This finding is particularly relevant as a cautionary tale for the systems neuroscience community modeling animal behaviors by drawing comparisons to RNNs (Mante et al., 2013) (Fig. 3). (iii) Utilizing analytical models, we demonstrate theoretical scaling laws and synthesize a phase diagram of mechanisms for solving the task, which is a function of the task properties and optimization parameters (Figs. S2 and 4G). (iv) Through a large-scale study of RNNs [^2] The dataset is publicly available through https://doi.org/10.5281/zenodo.15529757. To obtain this dataset, we used several computers with NVIDA RTX 3090 GPUs or equivalent, which roughly amounts to 230 kg CO2 emission, 930 km driven by an ICE car, 115 kg coal burned, 4 tree seedlings sequesting carbon for 10 year as computed via https://mlco2.github.io/impact/., we extract the same scaling laws in RNNs trained to perform delayed cue-discrimination tasks, which are quantitatively consistent with our theoretical explorations within error bars. We show that RNNs have little to no trouble learning a modified task with no delayed output, providing further evidence that training biologically inspired neural networks to delay their outputs is fundamentally difficult (Figs. 4, S3, S4, S5 and S6).
@@ 2 Background
Systems neuroscience often utilizes animal models, e.g., mice, rats, birds, and flies, to study cognitive functions including short-term memory (Serences, 2016). Alternatively, it is often possible to design artificial models such as RNNs, which can be trained on observed neural activities or tasks that are traditionally performed by animals in laboratories (Mante et al., 2013; Sussillo & Barak, 2013; Yang et al., 2019; Masse et al., 2019; Dubreuil et al., 2022). Then, once RNNs are trained, their computation can be reverse-engineered to investigate the nature of the neural computation performed by biological networks, generating hypotheses that can later be tested with experimental procedures (Walker et al., 2019; Finkelstein et al., 2021). Consequently, RNNs have recently become a key component of computational neuroscience studies (Dubreuil et al., 2022; Valente et al., 2022; Finkelstein et al., 2021; Sussillo & Barak, 2013).
Here, we consider a family of RNNs that are biologically relevant and whose parameters can be interpreted as functional connections between neurons (Perich & Rajan, 2020; Perich et al., 2021). Specifically, we study RNNs defined by the following equation:
@@
where $r(t)\in\mathbb{R}^{N}$ refers to the firing rates of $N$ neurons, $u(t)\in\mathbb{R}^{N_{\rm in}}$ pre-defined ${N_{\rm in}}$-dimensional inputs, $W\in\mathbb{R}^{N\times N},W^{\rm in}\in\mathbb{R}^{N\times{N_{\rm in}}}$ weight parameters with $b\in\mathbb{R}^{N}$ corresponding bias terms, and $\epsilon$ some noise that is, in practice, sampled i.i.d. from a Gaussian distribution. The output is taken as a projection of the neural activities as $\hat{o}(t)=f\left(W^{\rm out}r(t)+b_{\rm out}\right)$, where $f(\cdot)$ is either identity or sigmoid function in this work (see below) and $W^{\rm out}\in\mathbb{R}^{N_{\rm out}\times N_{\rm rec}}$ and $b_{\rm out}\in\mathbb{R}^{N_{\rm out}}$ for an $N_{\rm out}$-dimensional output. For notational simplicity, we assume $\epsilon=0$ until our empirical discussions in Section 3.4.
RNN models of sequential activity–Recent work has identified the generation of neural sequences as a potential key component of short-term memory (Rajan et al., 2016). This distributed sequence can be used to store the short-term memory content and explicitly keep track of time, which contrasts with earlier theories of short-term memory maintenance relying on the persistent activities of some neurons (Fuster & Alexander, 1971; Atkinson & Shiffrin, 1968) (Fig.1B). Previous work has provided insights into how these sequences are generated by examining neuronal-level mechanisms (Rajan et al., 2016; Sommer & Wennekers, 2005; Laje & Buonomano, 2013), often focusing on the learned connectivity structures (Rajan et al., 2016; Laje & Buonomano, 2013). However, the population-level causal mechanisms underlying neural computation, which could be investigated through latent dynamics, remain largely unexplored.
Figure 2: Changing optimization parameters may result in RNNs learning distinct attractor mechanisms. We trained a rank-2 RNN to perform the delayed activation task, in which RNNs initialized at a particular state (Methods) should delay their output by $T_{\rm delay}$ for a $T_{\rm resp}$ time interval. The task does not constrain the network activity or the output outside of these time windows. A Left. The normalized gradient norm and the loss function values. Middle. Flow maps of the learned latent dynamical systems demonstrating the (learned) attractor mechanism. The shaded areas around the decision boundary correspond to network outputs within $0.75$ (outer green boundary) and $0.25$ (outer red boundary). Right. Absolute firing rates of the network, whose latent dynamics were illustrated in the middle column. Dotted line corresponds to the start of the response period. Parameters: $N=100$ neurons, $\tau=10$ms, $\Delta t=5$ms, $T_{\rm delay}=75$ms, $T_{\rm resp}=10$ms, $\alpha=1$ for stochastic gradient descent using otherwise the default parameters in Pytorch (Paszke et al., 2017). B-C Same as in panel (A), but trained with learning rates of $10^{-2}$ and $10^{-4}$, respectively.
To study the latent mechanisms of sequential neural activities, we focus on a class of low-rank RNNs (i.e., when $W=\sum_{k=1}^{K}m^{(k)}n^{(k)T}$ are constrained to be low-rank). In this class, the computation performed by RNNs can be studied with the latent dynamical system (Dubreuil et al., 2022; Valente et al., 2022; Dinc et al., 2025a, 2023):
$\begin{split}\tau\dot{\kappa}_{k}(t)=&-\kappa_{k}(t)+\sum_{i=1}^{N}n_{i}^{(k)}\times\\ &\times\tanh\left(\sum_{l=1}^{K}m^{(l)}\kappa_{l}(t)+W^{\rm in}u(t)+b\right),\end{split}$ (2)
where the latent variables are defined as $\kappa_{k}(t)=n^{(k)T}r(t)$ (Dinc et al., 2025a). Intuitively, by constraining the rank of the weight matrix, $W$, one can project the computation performed by $N$ neurons down to a $K$-dimensional latent dynamical system. In this work, we use this paradigm (i.e., training low-rank RNNs) to illustrate the distinct mechanisms that support sequence generation, as they emerge during training. Then, our large-scale experiments generalize our findings to full-rank RNNs with no explicit restrictions.
@@ 3 Results
3.1 Latent mechanisms underlying neural sequences
@@
As a first step, we set out to answer our first question (i.e., Q1: What mechanisms support memory maintenance through sequential neural activity?) using a simple short-term memory task, stripped of any extraneous components, that can be solved using neural sequences (Fig. 1D-E). Specifically, we focused on one of the simplest tasks called a delayed activation task (Dinc et al., 2025b). In this task, the network is initialized to a particular state, which is randomly selected in the state space (Methods), and has to withhold its response ($\hat{o}(t)=0$) for a $T_{\rm delay}$ time window. Afterwards, the network is forced to output a response ($\hat{o}(t)=1$) for $T_{\rm resp}$ time window.
Two distinct mechanisms can subserve neural sequences–Broadly, we set out to study two latent mechanisms that may subserve the observed neural sequences. The first mechanism relies on the generation of a set of slow-points, neural states in which the neural dynamics vary slowly ($\dot{r}(t)\approx 0$). In this mechanism, neural sequences emerge during the slow transition of the latent dynamical system through these slow-points. This transition is often structured, i.e., the same latent trajectory would be followed as the system passes through these regions (Sussillo & Barak, 2013), and therefore can lead to stable and sequential activations of neurons (Fig. 1B). The second mechanism relies on the generation of limit cycles, which are defined as closed trajectories that represent sustained, periodic oscillations. In this mechanism, when the oscillation periods are much larger than the trial windows, neural sequences can emerge during the task execution without repetition (Fig. 1C). But, once the task concludes, the limit cycles would continue to oscillate and regenerate the sequential activities, whereas the slow-points are transitionary and would not repeat (Fig. 1B-C).
Since a minimum of two dimensions is required to generate oscillatory behavior in dynamical systems (Strogatz, 2018), we start by studying the learned latent representations in an RNN with rank $K=2$ (corresponding to two-dimensional latent dynamics). The output, $\hat{o}(t)$, of the RNN is defined as a nonlinear projection of the latent variables using a sigmoid function $f(\cdot)$. For the largest learning rate (Fig. 2A), this rank-2 RNN was incapable of solving the task despite the nearly-zero gradient [^3] This phenomenon has been studied in details by a previous work (Dinc et al., 2025b), in which the learning dynamics can get stuck in no-learning zones with near-zero gradients when very large learning rates are used.. As the learning rates decreased, it was capable of solving the task with limit cycles with a moderate learning rate (Fig. 2B), but utilized slow-points (also referred to as “SP manifolds” for generality) for the lowest learning rate (Fig. 2C). Notably, both mechanisms produced neural activity sequences during the task window (Fig. 2, right panels).
Figure 3: Changing trivial task parameters may result in RNNs favoring limit cycle solutions. A We performed experiments similar to those in Fig. 2, but now with an added post-reaction period ($T_{\rm post}$). B Learned latent dynamics for the fully trained networks in panel (A). Parameters for A-B: As in Fig. 2A, with $T_{\rm delay}=30$ms, $T_{\rm resp}=10$ms, $T_{\rm post}=0\text{ or }30$ms, and $\alpha=10^{-3}$. C Fraction of networks developing rotating solutions (i.e., limit cycles) when solving the two versions of the task as a function of delay $T_{\rm delay}$. Parameters: $\alpha=0.005$, varying levels training epochs corresponding to $T_{\rm delay}$, otherwise the same as in panels A-B. Comparisons: Fisher’s exact test with Bonferroni corrections, ${}^{\ast\ast\ast}p<0.001$, and ${}^{\ast\ast}p<0.01$.
@@
Despite leading to equivalent neural activities within the trial window, the two mechanisms had qualitatively different behaviors after the trial concluded (Fig. 2). Limit cycles had led to recurrent generations of the same sequences, whereas the SP manifolds converged to a persistent activity style representation after the trial concludes. We will make this observation rigorous in our large-scale studies below. For now, we provide insights into another relevant question: Do RNNs commit to learning one mechanism from the start, or are they capable of changing their inner mechanisms during the training?
@@
Learned latent mechanisms can evolve during learning–We next studied the latent dynamical systems of a rank-2 RNN as it was training to solve the delayed activation task (Fig. S1). Interestingly, the loss function showed characteristic periods of learning, corresponding to changes in mechanisms (Fig. S1A). Specifically, after the first jump in the loss function, a SP manifold was formed in the latent dynamical system (Fig. S1B; left). Here, a line of slow-points allowed the network output to stall until $T_{\rm delay}$. Then, the SP manifold had acquired a curved structure (Fig. S1B; middle), reminiscing an onset to a rotating solution. Finally, after the oscillations in the loss function values ceased, a limit cycle had emerged in the latent dynamical system (Fig. S1B; right). Hence, we confirmed that RNNs could change their inner mechanisms during learning, i.e., what started as a SP manifold solution later evolved into a limit cycle.
@@
In reality, since functional connections in the brain can rapidly reorganize within the start and end of a trial (Ebrahimi et al., 2022), the two mechanisms can become practically inseparable in biological neural networks. In other words, there is an equivalence class of mechanisms that can generate neural sequences that have the same local behavior within the task window, but can show distinct properties outside; and the two mechanisms we identified belong to this class. For the rest of this work, we will formalize this connection between the learned latent mechanisms and the task and/or optimization parameters. In doing so, we will generalize these findings beyond anecdotal observations via theoretical investigations and large-scale experiments.
3.2 Influence of task design on the learned mechanism